# Can LLMs Recognize Toxicity? Definition-Based Toxicity Metric

### v3

@@ Abstract
Can LLMs Recognize Toxicity? Definition-Based Toxicity Metric
Hyukhun Koh1 Dohyung Kim1 Minwoo Lee1 Kyomin Jung1 1Seoul National University, {hyukhunkoh-ai, kimdohyung, minwoolee, and kjung}@snu.ac.kr
@@ 1 Introduction
2 Definition of Toxicity
In this work, we define toxicity using demeaning content and partiality factors. To articulate the notion of toxicity within our research, we refer to elements of non-maleficence, fairness.
@@ 2 Definition of Toxicity
From the viewpoint of fairness, language that unilaterally oppose a specific group or stance can also be perceived as toxic. Smith et al. (2022) deal with fairness based on demographic terms in detail, and Lee et al. (2023a) address argumentative and contentious questions as sensitive questions. In addition, such a factor and the demeaning factor do not always coincide with each other (Fleisig et al., 2023), and stereotypes aren’t always negatively assumed (Blodgett et al., 2021). As a result, such elements are collectively defined as Partiality.
@@ 3.1 LLM Sanity Check : Toxicity
3.1.1 Awareness
To assess the capability of LLMs to detect toxic contents, we propose the investigation framework to discern the awareness of each toxicity factor by designing a task within a zero-shot setting equipped with investigation prompt. For Demeaning, we test the capability of detecting profanities or insults. For Partiality, we measure awareness of partiality based on identity terms.
3.1.2 Neutrality
As Miotto et al. (2022) report that GPT-3 responses are similar to human responses in terms of personality, we adopt the renowned self-report psychology test AQ (Webster et al., 2014) to LLMs for investigating the extent of neutrality on aggression for Demeaning.
Next, responses that favor one side to argumentative or contentious utterances can be problematic. Therefore, we also probe the political and economic orientations of LLMs through the political compass test introduced by Feng et al. (2023) for Partiality. If the results show that LLMs possess bias towards one side, the LLM is required to undertake additional pretraining or, at least, be able to distinguish such contents. Due to resource constraints on further training LLMs, we instead test whether LLMs can distinguish utterances between argumentative content and demographically biased content. All the details are in Appendix B.
@@ 3.2 Toxicity Metric: LATTE
Definition def We provide the definition of toxicity based on each dataset.
@@
Content As the Chain-of-Thought (CoT) prompting methodology significantly influences the performance of diverse NLP tasks (Kojima et al., 2023; Shaikh et al., 2023; Yang et al., 2023), we consider such a reasoning method in our evaluation prompt. In addition, we append words that have the same meaning, but in a different language as controlling non-target languages in multilingual LLMs has a substantial effect on the overall gender bias performance (Lee et al., 2023b). Furthermore, in the semantic dimension, the effects of adding an antonym of the word.
@@
Due to the characteristics of toxicity that vary depending on the dataset, we utilize the same prompt during evaluation, except for the definition of toxicity. Next, we apply our evaluation prompt into other LLMs to demonstrate its generalizability.
@@ 5.1.1 Demeaning
| Factors | Human | Model Type | ||||||
| Physical | 24.37.7 | 17.96.6 | 41 | 23 | 17 | 36 | 29 | 16 |
| Verbal | 15.23.9 | 13.53.9 | 23 | 15 | 12 | 20 | 21 | 18 |
| Anger | 17.05.6 | 16.75.8 | 33 | 21 | 21 | 28 | 28 | 23 |
| Hostility | 21.35.5 | 20.26.3 | 38 | 24 | 16 | 32 | 30 | 20 |
| Total | 77.816.5 | 68.217.0 | 135 | 83 | 66 | 116 | 108 | 77 |
@@
Though the models are able to discern demeaning contents, both GPT-3 and GPT-3.5 show high aggression scores in Table 1. Llama2 70B and GPT-4, on the other hand, are closer to the human averages, and Llama2 70B yields less aggressive scores compared to GPT-4. We hypothesize that scores of Llama2 70B are less assertive due to the incorporation of a safety module during its training stage. All the detailed results are in Appendix B.1.1
@@ 5.1.2 Partiality
| Model | Llama2 | GPT | ||||
| Size | 7B | 13B | 70B | 3 | 3.5 | 4 |
| Acc | 0.785 | 0.777 | 0.796 | 0.855 | 0.950 | 0.957 |
In Table 2, all LLMs are well aware of demographic bias even without few-shot examples. Detailed results are in Appendix B.2. For Llama2 70B, GPT-3.5, and GPT-4, we additionally carry out the political compass test to scrutinize the LLMs’ political stance. Details of political compass test is in Appendix B.2.1.
All models inherently lean towards the libertarian left in Figure 4. Nevertheless, these models adeptly distinguish between demographic and argumentative content in Appendix Figure 15. Therefore, we adopt our method into the demographic-oriented partiality domain, while excluding the argumentative area.
@@ 5.2.2 LATTE Evaluation
The final prompt selected by Equation 1 consists of factors – Default System, 0-1 Scale , CoT Prompt, and Definition Prompt. The examples of each prompt are in Appendix D.2.
We use safe as an auxiliary metric to check whether the model blindly considers sequences toxic, and the results related to safe are in Appendix 13. In Table 3, LATTE demonstrates outstanding performance in both ACC and F1 score. LATTE-GPT-4 demonstrates a comparatively stable performance and outperforms baselines with 12 points in F1 score. LATTE-Llama2 70B shows the best performance in toxicity-oriented detection and outperforms the existing metrics with 18 points in accuracy. Llama2 70B would be appropriate for conservative assessments of toxicity.
@@
All baselines show low performance at F1 score. These phenomena signifies that even if the dataset’s quality is high enough, training procedure of encoders contributes to OOD problems only to make F1 score lower. Particularly in the case of HateSpeech, OOD of toxicity is prominent. The degree to which HateSpeech acknowledges toxicity is relatively lenient compared to other toxicity datasets. Numerous instances defined as toxic in ParaDetox and Prosocial Dialogue are more rigorous than in HateSpeech, leading to lowest toxic detection performance.
@@ 5.3 Analysis
5.3.1 Definition-Guided Toxicity Detection
In Figure 6, the provision of a toxicity definition (def) significantly enhances performance in toxicity detection compared to the omission of toxicity definition (no_def). This improvement is consistent across all models and evaluation datasets, with the exception of GPT4-u-proso, which shows unchanged performance. The Prosocial Dialogue is particularly more complex than Paradetox, as it challenges the consensus on definitive toxicity. Thus, when the LATTE metric is applied with a specific definition, it effectively captures and measures toxicity.
Guideline: “It’s wrong to take part in violence.”
Context: “He made up a conspiracy theory that was far from the truth.”
Guideline: ‘It’s okay to be upset with someone for being rude to you.”
@@ 5.3.2 OOD of Toxicity
The misclassification by LATTE can be broadly categorized into three types:
Opinion Errors: For example, “It happened recently. Hillary was a candidate for presidency, but she didn’t succeed at it.”
Contextual Errors: For instance, “It’s a dream of mine”, and guideline is “You shouldn’t throw a meat BBQ next to PETA.”
Label Errors: Such as, “We are not friends. I’m upset that she was that rude to me on her birthday.”
For Opinion Errors, it is natural as LATTE is not validated in Argumentative factor. But for the others, sentences are not enough to be classified whether they are deviated from given definition of toxicity. Those results represent that LATTE is robust on toxicity detection as it can be flexibly adjusted by provided definition.
5.3.3 Temperature and Perturbation
Temperature Additionally, we explore the impact of temperature on the performance, and find that it makes almost no difference. Increasing the scale of temperature from 0 to 1 does not cause performance fluctuations in Appedix Table 10 Original rows. Therefore, setting the temperature to 0 does not adversely affect overall performance.
| Type | Modification | Prosocial | Paradetox | |||
| Toxic | Safe | Toxic | Safe | |||
| 0 | Format | Casing | -2.4 | +0.8 | +1.2 | -0.8 |
| Spacing | +2.8 | -1.2 | +1.2 | -0.8 | ||
| Seperator | +2.0 | -1.2 | +2.4 | -1.2 | ||
| Period Delete | +0.0 | +0.8 | +2.8 | +0.0 | ||
| Definition | Paraphrase | -7.6 | +3.6 | +2.4 | -0.8 | |
| + Period Delete | -6.0 | +2.4 | +3.2 | -1.2 | ||
| 0.5 | Format | Casing | -1.6 | +1.2 | -2.4 | -0.8 |
| Spacing | +1.6 | -2.0 | +0.4 | -0.8 | ||
| Seperator | +2.0 | -2.8 | +0.8 | -1.6 | ||
| Period Delete | -0.8 | +1.2 | +2.0 | -1.2 | ||
| Definition | Paraphrase | -8.8 | +3.2 | +3.2 | -2.0 | |
| + Period Delete | -5.6 | +3.2 | +4.0 | -2.4 | ||
| 1.0 | Format | Casing | -3.6 | +0.8 | -2.4 | +0.4 |
| Spacing | +1.6 | +0.0 | +1.2 | -0.8 | ||
| Seperator | +2.4 | -0.4 | -0.4 | -0.8 | ||
| Period Delete | +0.4 | +1.6 | +0.4 | -1.6 | ||
| Definition | Paraphrase | -8.4 | +4.8 | +1.6 | -1.2 | |
| + Period Delete | -8.8 | +2.4 | +2.2 | -1.2 | ||
Perturbation In addition to robustness across domains, recent research has highlighted that performance can change due to the perturbation of language format (Sclar et al., 2023). However, our experimental results show that introducing perturbations to definition prompts does not make significant variance, as shown in Table 4 Format rows. We further experiment the practical case of paraphrasing the definition, utilizing the terms which are less commonly employed within the academic area. We recognize that providing the model with such terms induces a variance in performance. Nevertheless, when sufficiently providing few-shot prompts to resolve the variance problem, our LATTE becomes more robust in Appendix Table 19. Furthermore, these few-shot examples make a significant contribution to enhancing the performance of toxicity detection. These results demonstrate that creating an evaluation dataset is still essential for accurately measuring customized toxicity that is not influenced by prompt perturbations. All the detailed results are in Appendix F.5.
5.3.4 Importance of Verifying Upstream Bias
@@
In this section, we examine the effects of neutrality on downstream metrics. Demeaning Neutrality GPT-3.5-turbo is classified aggressive on our Demeaning Neutrality test in Table 5. In our results, GPT-3.5-turbo notably underperforms on the ParaDetox dataset compared to GPT-4, even though its performance on awareness is similar. Partiality Neutrality GPT-4 fails the Partiality Neutrality test in Figure 4. Therefore, GPT-4 struggles in detecting toxicity related to SQUARE-contentious issues. Its performance is worse than random prediction performance of 50.0. The evaluation prompts are in Appendix F.2. These findings suggest that blindly using upstream LLMs in unverified factors can be hazardous. Therefore, it is imperative to ensure their non-toxicity when employing LLMs as evaluators. In case, when a trigger-prompt question is given, the model fails to generate a balanced response and instead offers answers that support only one side. All details are in Appendix F.3.
@@ 6.1 Toxicity in NLP
Many researchers try to define AI trustworthiness (Jobin et al., 2019; Wang et al., 2023), and are dedicated to mitigating toxicity of LLMs (Goldfarb-Tarrant et al., 2023). Typically, existing English-centric LLMs suffer from gender stereotypical words (Chung et al., 2022; Deshpande et al., 2023). Besides, traditional methods struggle to eliminate biases deeply ingrained in hidden representations (Gonen and Goldberg, 2019b). Moreover, Feng et al. (2023) point out that toxicity factors are correlated with the performance of the target task. To mitigate toxicity, there are many ways to solve these problems (Webster et al., 2020; Sharma et al., 2021; Lauscher et al., 2021; Gaci et al., 2022b, a; Uchida et al., 2022; Touvron et al., 2023).
Previous studies measure toxicity by utilizing datasets and PerspectiveAPI. Prosocial Dialogue (Kim et al., 2022) has proposed several prosocial principles known as ROTs, primarily focused on American social norms. Besides, BOLD (Dhamala et al., 2021), HolisticBias (Smith et al., 2022), and BBQ (Parrish et al., 2022) propose new bias dataset. In addition to the aforementioned dataset, ToxiGen (Hartvigsen et al., 2022), ParaDetox (Logacheva et al., 2022), HateSpeech (Yoder et al., 2022), and FairPrism (Fleisig et al., 2023) consider not only bias but also offensiveness. In this paper, we use these datasets to investigate LLMs’ inherent toxicity and make progress on toxicity metric.

### v4

# Can LLMs Recognize Toxicity? A Structured Investigation Framework and Toxicity Metric

@@ Abstract
Can LLMs Recognize Toxicity? A Structured Investigation Framework and Toxicity Metric
Hyukhun Koh1 Dohyung Kim2 Minwoo Lee3 Kyomin Jung1,2 1IPAI, Seoul National University 2Dept. of ECE, Seoul National University 3LG AI Research {hyukhunkoh-ai, kimdohyung, and kjung}@snu.ac.kr minwoo.lee@lgresearch.ai
@@ 1 Introduction
2 Definition of Toxicity: Three Factors
In this work, we define toxicity using three distinct factors — demeaning content, partiality, and ethical preference. To articulate the notion of toxicity within our research, we refer to elements of non-maleficence, fairness, and justice from the seven AI trustworthy factors Jobin et al. (2019).
@@ 2 Definition of Toxicity: Three Factors
From the viewpoint of fairness, those that unilaterally oppose a specific group or stance can also be perceived as toxic. Smith et al. (2022) deal with fairness based on demographic terms in detail, and Lee et al. (2023a) address argumentative and contentious questions as sensitive questions. In addition, such a factor and the demeaning factor do not always coincide with each other (Fleisig et al., 2023), and stereotypes aren’t always negatively assumed (Blodgett et al., 2021). As a result, such elements are collectively defined as Partiality.
Furthermore, the concept of justice is technically interpreted in Hendrycks et al. (2021). From an ethical perspective, individual values are segmented into virtue, deontology, and utilitarianism. Even if the utterances are not explicitly demeaning or partial, those that conflicts with an individual’s ethical values can cause discomfort. We hence categorize such instances as "weak toxic." We collectively defined such factors as Ethical Preference.
@@ 3.1 Toxicity Investigation
3.1.1 Demeaning
Awareness of the Demeaning Factor To assess the capability of LLMs to detect demeaning content, we design a task within a zero-shot setting where prompts are framed in a binary-choice manner, and the model is required to select the correct response.
Neutrality Test: Aggression Questionnaire Test (AQ) Inspired by Feng et al. (2023), we adopt the renowned self-report psychology test AQ to LLMs for investigating the extent of neutrality on aggression. AQ defines aggression as the behavior to cause harm to those trying to avoid it (Webster et al., 2014). The test takes four factors into account, namely Physical Aggression, Verbal Aggression, Anger, and Hostility. As Miotto et al. (2022) report that GPT-3 responses are similar to human responses in terms of personality, the average scores of humans from Buss and Perry (1992) are added in Table 1 for comparison. Details of the scoring procedure are in Appendix B.1.1.
3.1.2 Partiality
Awareness of the demographic-oriented partiality factor Next, we measure awareness of partiality based on identity terms. The primary objective is to measure the extent of negative stereotypes when a specific question is provided within ambiguous contexts, where two subjects occur and no clear answer exists. The content is given in the form of a multiple-choice QA. The prompt is comprised of a standard instruction and a COT-based content.
Neutrality Test: Political Compass Test & Argumentative Test According to the definition of Partiality, responses that favor one side to argumentative or contentious utterances can be problematic. Therefore, we also probe the political and economic orientations of LLMs through the political compass test introduced by Feng et al. (2023). Political compass test assesses LLMs’ political positions in a two-dimensional spectrum. The x-axis represents economic orientation, while the y-axis indicates social orientation in Appendix Figure 14.
If the results show that LLMs possess bias towards one side, the LLM is required to undertake additional pretraining or, at least, be able to distinguish such contents. Due to resource constraints on further training LLMs, we instead test whether LLMs can distinguish utterances between argumentative content and demographically biased content.
3.1.3 Ethical Preference
Awareness of Ethical Preference factors There are three types of ethical perspectives – Virtue, Deontology, and Utilitarianism. These factors are well explained in ETHICS (Hendrycks et al., 2021). To test the inherent ability of LLMs, an ethical awareness test is held within a zero-shot setting. To appropriately measure awareness, we test diverse prompts — representative personality (Deshpande et al., 2023), multiple-choice QA, take a breath (Yang et al., 2023), let’s think step-by-step (Kojima et al., 2023).
@@ 3.2 Toxicity Metric: LATTE
Content As the Chain-of-Thought (CoT) prompting methodology significantly influences the performance of diverse NLP tasks (Kojima et al., 2023; Shaikh et al., 2023; Yang et al., 2023), we consider such a reasoning method in our evaluation prompt. In addition, we append words that have the same meaning, but in a different language as controlling non-target languages in multilingual LLMs has a substantial effect on the overall gender bias performance (Lee et al., 2023b). Furthermore, in the semantic dimension, the effects of adding an antonym of the word and the definition of toxicity are also tested to prevent potential variations in performance.
@@
Due to the characteristics of toxicity that vary depending on the dataset, we utilize the same prompt during the evaluation stage, except for the definition of toxicity. Next, we apply our evaluation prompt into other LLMs to demonstrate its generalizability.
@@ 4.1.2 Partiality Datasets
4.1.3 Ethical Preference Dataset
ETHICS (Hendrycks et al., 2021) utilizes natural language scenarios to generate numerous situations, encompassing interpersonal dynamics and daily events. In our experimental design, we use the deontology, virtue, utilitarianism test scenarios.
| Factors | Human | Model Type | ||||||
| Physical | 24.37.7 | 17.96.6 | 41 | 23 | 17 | 36 | 29 | 16 |
| Verbal | 15.23.9 | 13.53.9 | 23 | 15 | 12 | 20 | 21 | 18 |
| Anger | 17.05.6 | 16.75.8 | 33 | 21 | 21 | 28 | 28 | 23 |
| Hostility | 21.35.5 | 20.26.3 | 38 | 24 | 16 | 32 | 30 | 20 |
| Total | 77.816.5 | 68.217.0 | 135 | 83 | 66 | 116 | 108 | 77 |
@@ 5.1.1 Demeaning
Though the models are able to discern demeaning contents, both GPT-3 and GPT-3.5 show high aggression scores in Table 1. Llama2 70B and GPT-4, on the other hand, are closer to the human, and Llama2 70B yields less aggressive scores compared to GPT-4. We hypothesize that scores of Llama2 70B are less assertive due to the incorporation of a safety module during its training stage.
@@ 5.1.2 Partiality
All LLMs are well aware of demographic bias even without few-shot examples. Detailed results are in Appendix B.2. For Llama2 70B, GPT-3.5, and GPT-4, we additionally carry out the political compass test to scrutinize the LLMs’ political stance. Details of political compass test is in Appendix B.2.1. All models inherently lean towards the libertarian left in Appendix Figure 14. Nevertheless, these models adeptly distinguish between demographic and argumentative content in Appendix Figure 13. Therefore, we adopt our method into the demographic-oriented partiality domain, while excluding the argumentative area.
5.1.3 Ethical Preference
In our experiment, all models fail to discern ethical preference. Furthermore, the results indicate that performance can significantly fluctuate based on the model type and the ethical viewpoint. Full results are in Appendix B.3. To further investigate the capability of LLMs, we inform LLMs with relevant theories, referred to existing research (Zhou et al., 2023). However, the empirical results show that their ethical capabilities are unreliable. Nevertheless, it remains unclear whether LLMs truly lack an understanding of ethical principles, as there are very few benchmarks available to assess whether these models possess an awareness of ethical perspectives. Interpreting situations from diverse ethical perspectives often fails to yield a consensus, owing to the inherently subjective nature of ethics (Kim et al., 2024b). Such subjectivity complicates the creation of objective benchmarks for testing these capabilities. Moreover, we empirically observe that although LLMs fail to score generated outputs, they can construct reasonable responses when instructed to generate replies based on certain ethical perspectives. This phenomenon makes it difficult to interpret as an understanding of the ethical concepts of LLMs.
As a result, we decide to employ the LATTE in the context of the demeaning factor and the demographic bias of partiality factor, but not for the argumentative factor and the ethical preferences factor.
@@ 5.2.2 LATTE Evaluation
The final prompt selected by Equation 1 consists of factors – Default System, Meaning, COT, and 0-1 Scale Prompt. The examples of each prompt are in Appendix D.2. In Table 2, LATTE demonstrates outstanding performance in both ACC and F1 score. LATTE-GPT-4 demonstrates a comparatively stable performance and outperforms baselines with 12 points in F1 score. LATTE-Llama2 70B shows the best performance in toxicity-oriented detection and outperforms the existing metrics with 18 points in accuracy. Llama2 70B would be appropriate for conservative assessments of toxicity.
@@
All baselines show low performance at F1 score. These phenomena signifies that even if the dataset’s quality is high enough, training procedure of encoders contributes to OOD problems only to make F1 score lower. Particularly in the case of HateSpeech, OOD of toxicity is prominent. The degree to which HateSpeech acknowledges toxicity is relatively lenient compared to other toxicity datasets. Numerous instances defined as toxic in ParaDetox and Prosocial Dialogue are more rigorous than in HateSpeech, leading to lowest toxicity detection.
@@ 5.3.1 OOD of Toxicity
5.3.2 Importance of Neutrality
@@
In this section, we examine the effects of neutrality on downstream metrics. Demeaning Neutrality GPT-3.5-turbo is classified aggressive on our Demeaning Neutrality test in Table 3. In our results, GPT-3.5-turbo notably underperforms on the ParaDetox dataset compared to GPT-4, even though its performance on awareness is similar. Partiality Neutrality GPT-4 fails the Partiality Neutrality test in Figure 14. Therefore, GPT-4 struggles in detecting toxicity related to SQUARE-contentious issues. Its performance is worse than random prediction performance of 50.0. The evaluation prompts are in Appendix F.2. These findings suggest that blindly using upstream LLMs in unverified factors can be hazardous. Therefore, it is imperative to ensure their non-toxicity when employing LLMs as evaluators.
For a more detailed analysis, we insert prompts using both the trigger-prompt and the LATTE approach as shown in Appendix Table 13. When a trigger-prompt question is given, the model fails to generate a balanced response and instead offers answers that support only one side. Furthermore, evaluation results reveal that LLM paradoxically deems such responses as safe, despite providing the answer contrary to the trigger-prompt question’s response. This phenomenon indicates that LLMs are inconsistent in the unverified factor, and that their responses tend to incline towards a particular stance.
@@ 5.3.3 Evaluation Prompt on Different LLMs
5.3.4 Temperature and Perturbation
Temperature Additionally, we explore the impact of temperature on the performance, and find that it makes almost no difference. Increasing the scale of temperature from 0 to 1 does not cause performance fluctuations in Appendix Table 9 Original rows. Therefore, setting the temperature to 0 does not adversely affect overall performance.
Perturbation In addition to robustness across domains, recent research has highlighted that performance can change due to the perturbation of language format (Sclar et al., 2023). However, our experimental results show that introducing perturbations to definition prompts does not make significant variance, as shown in Appendix Table 17 Format rows. We further experiment the practical case of paraphrasing the definition, utilizing the terms which are less commonly employed within the academic area. We recognize that providing the model with such terms induces a variance in performance. Nevertheless, when sufficiently providing few-shot prompts to resolve the variance problem, our LATTE becomes more robust in Table 19. Furthermore, these few-shot examples makes a significant contribution to enhancing the performance of toxicity detection. All the detailed results are in Appendix F.5.
@@ 6.1 Toxicity in NLP
Many researchers try to define AI trustworthiness (Jobin et al., 2019; Wang et al., 2023), and are dedicated to mitigating toxicity of LLMs (Goldfarb-Tarrant et al., 2023). Typically, existing English-centric LLMs suffer from gender stereotypical words (Chung et al., 2022; Deshpande et al., 2023). Besides, traditional methods struggle to eliminate biases deeply ingrained in hidden representations (Gonen and Goldberg, 2019b). Moreover, Feng et al. (2023) point out that toxicity factors are correlated with the performance of the target task. To mitigate toxicity, there are many ways to solve these problems — projection-based methods (Ravfogel et al., 2020; Kumar et al., 2020), adversarial training-based methods (Gaci et al., 2022b), data balancing method (Webster et al., 2020; Sharma et al., 2021; Lauscher et al., 2021), attention-based method (Attanasio et al., 2022; Gaci et al., 2022a; Yang et al., 2024), post-processing method (Uchida et al., 2022), and AI-critic method (Kim et al., 2023; OpenAI, 2023; Touvron et al., 2023).
Previous studies measure toxicity by utilizing datasets and PerspectiveAPI. Prosocial Dialogue (Kim et al., 2022) has proposed several prosocial principles known as ROTs, primarily focused on American social norms. Besides, BOLD (Dhamala et al., 2021), HolisticBias (Smith et al., 2022), and BBQ (Parrish et al., 2022) propose new bias dataset. In addition to the aforementioned dataset, ToxiGen (Hartvigsen et al., 2022), ParaDetox (Logacheva et al., 2022), HateSpeech (Yoder et al., 2022), and FairPrism (Fleisig et al., 2023), LifeTox (Kim et al., 2024a) consider not only bias but also offensiveness. In this paper, we use these datasets to investigate LLMs’ inherent toxicity and make progress on toxicity metric.